# Bone: Block Affine Transformation as Parameter Efficient Fine-tuning Methods for Large Language Models

### v2

@@ Abstract
Low-Rank Adaptation (LoRA) has achieved remarkable training results by freezing the original weights and training only low-rank matrices, establishing itself as the predominant fine-tuning method for LLMs. In pursuit of performance closer to full-parameter training, a series of LoRA variants have emerged, such as LoRA+, PISSA, Olora, and LoRA-GA. However, these improvements complicate the initial setup of model training and increase initialization time. More importantly, they overlook the internal interactions of the original weight information. To address these issues, we introduce a novel theory, “Weight Guide” aimed at continuously guiding trainable matrices through the original weights during training to enhance the utilization of weight information. Based on this theory, we designed a new PEFT technique called Bone (Block Affine), which not only enhances the utilization of original weight information but also emphasizes the internal connections between weights, leading to faster convergence and better data fitting. Experimental comparisons across two different LLM architectures (LLaMA2, RWKV6) and various parameter scales demonstrate that the Bone structure can achieve rapid convergence and superior data fitting without the need for complex initialization. For example, when fine-tuning LLaMA2-7B on the MetaMathQA dataset and validating on GSM8k and math benchmarks, Bone achieved fine-tuning scores of 49.36 and 8.8, respectively, outperforming PISSA by 5.84% and 1.96%.
@@ 1 Introduction
LoRA significantly reduces memory usage by freezing the original weights $W$ and updating two low-rank matrices $A$ and $B$. Typically, either $A$ or $B$ is initialized to zero, ensuring that the initial state of LoRA is consistent with the pre-trained model, The Figure 2(a) illustrates the structure visualization. However, extensive experiments Ding et al. (2023); Liu et al. (2024b); Biderman et al. (2024) have shown that LoRA’s convergence is significantly slower compared to full fine-tuning. This slow convergence is likely due to the small gradients caused by the zero initialization of either $A$ or $B$. To address this issue, researchers have proposed several LoRA variants, such as LoRA+ Hayou et al. (2024), PISSA Meng et al. (2024), and LoRA-GAWang et al. (2024). Despite their excellent performance, these LoRA variants inevitably introduce complexity into the fine-tuning process and reduce structural flexibility. For example, LoRA+ requires manual adjustment of different learning rates for $A$ and $B$; PISSA necessitates performing SVD decomposition on $W$ at the beginning of training, which can be time-consuming when the model parameters are large.
In this paper, we propose a novel PEFT method called Bone (Block Affine) to overcome the limitations of LoRA variants. It is well-known that LLMs have extremely large parameters, with complex interactions among their internal weights, but most LoRA variants utilize original weight information minimally, with only a few leveraging it during initialization to enhance performance. To increase the utilization of this information and promote internal interactions among weights, we propose a novel concept called “Weight Guide” This approach ensures that the trainable matrices are consistently guided and constrained by the original weights throughout every step of the training process. As illustrated in Figure  1(b), LoRA variants interact with the original weights only once during initialization to extract essential components. In contrast, the Bone structure is guided (constrained) by the original weights after each update, significantly increasing the new weights’ utilization of the original weights. As the training steps increase, the influence of the original weights becomes more pronounced. This is evident from the loss curves, which show that the differences between Bone and other methods grow larger in the later stages of training.
To promote the internal interactions of information within the original weight matrix $W$, Bone employs an trainable zero-initiated matrix, named bone, to perform Block Affine operations on $W$. The specific calculations of Block Affine are detailed in Section 3, and the structural visualization is shown in Figure  2. To verify the feasibility of Bone, we conducted experiments on two different LLM architectures (LLaMA2 Xu et al. (2023), RWKV6 Peng et al. (2024)). As shown in Figure  3, Bone achieves the fastest convergence, and the results on the test set demonstrate its superior data fitting and generalization capabilities (Table  2, Table  3).
@@
1. We propose a new concept, “Weight Guide” which allows the continuous utilization of original weight information throughout the training process.
2. We propose a novel PEFT method called Bone (Block Affine) to overcome the limitations of LoRA variants. Block Affine not only efficiently promotes internal interactions of information within the weights but also enhances the weight utilization of low-rank matrices.
3. The effectiveness of the Bone structure and “Weight Guide” has been validated through extensive experiments.
@@ 2 Related Works
3 Bone: Block Affine
In this section, we will introduce the Bone (Block Affine) method. Our design is based on multiple considerations. First, to achieve the aforementioned “Weight Guide” We modify the formula $Y=X(W+\Delta W)$ by replacing $\Delta W$ with $W\Delta W+\Delta W$, allowing the original weights to continuously guide (constrain) the direction of $\Delta W$ throughout the training process. To verify whether the guidance (constraint) from the original weights has a positive effect on the trainable matrix, we conducted an ablation experiment in Section  4.5, we named the comparison structure “Bone-Unconstrained”, with the comparison formula: $\Delta W=WA+B$. The experimental results indicate that “Weight Guide” not only provides positive guidance but also enhances the utilization of matrix parameters.
Second, to enhance the fusion of internal weight features and extract more effective information, we opted for matrix multiplication over the Hadamard product in the $W\Delta W$ operation. The Hadamard product only captures features at corresponding positions, whereas matrix multiplication allows the accumulation of features from an entire row into a single feature. To validate this approach, we conducted experiments in Section  4.5 to compare different computational methods in terms of feature fusion and extraction. Here, we named the comparison structure “Bone-Hadamard” to facilitate readers’ understanding and reference in the following sections. Ultimately, we chose matrix multiplication as the core operation.
Finally, we hypothesize that the matrix $W$ has dimensions $(n,n)$, ideally, we train with a full-rank matrix $bone$ $(n,n)$. The specific calculation for Block Affine is as follows:
$Y=X(W+\Delta W)$ (1)
$\Delta W_{n,n}=W_{n,n}bone_{n,n}+bone_{n,n}$ (2)
We know that in LLMs, the matrix $W$ is typically very large, and in most cases, the number of rows and columns is not equal. Setting $bone$ to the same dimensions as W would result in excessive computational overhead, which goes against the principle of parameter-efficiency. Therefore, we adjust $bone$ to be a low-rank matrix. When the weight matrix $W$ has dimensions $(n,m)$, it is divided according to a block_size $b$ into $W(n/b,b,m/b,b)$, yielding $(n/b)\times(m/b)$ blocks. Inspired by GQA(Grouped Query Attention)Ainslie et al. (2023), we group these blocks together to significantly reduce the number of trainable parameters without affecting performance. Different groups share a single Bone matrix for computation. Thus, the Bone matrix is set to $(g,b,b)$, where $g$ is the group size. The $W$ matrix is reshaped to $(gn,g,b,b)$, where $gn$ represents the number of groups. The Bone structure is illustrated in Figure 2(b). The final formula can be expressed as:
$W_{gn,g,b,b}=Reshape(W_{gn*b,g*b})=Reshape(W_{n,m})$ (3)
$\Delta W_{n,m}=Reshape(W_{gn,g,b,b}bone_{g,b,b}+bone_{g,b,b})$ (4)
During our comparative experiments between RWKV6 and LLaMA2, we found that the matrix $W$ has several different dimensions. For example, in RWKV6-7B, the dimension of ffn’s $k$ is $(4096*3.5,4096)$, while in LLaMA2-7B, the dimension of gate_proj in the mlp layer is $(4096,11008)$. These differences require us to continuously adjust g to ensure the structure functions correctly. To better compare Bone with various LoRA variants and to address the compatibility issues arising from differences in FFN layers across different models and parameter settings, We set $g=n/b$ in $bone_{g,b,b}$ allowing the parameters of Bone to be equivalent to those of $LoraA_{n,r}$. During computation, Bone only requires a dimension transformation, The Bone structure is illustrated in Figure 2(c), and the modified formula is as follows:
$W_{m/b,n/b,b,b}=Reshape(W_{n,m})$ (5)
$bone_{n/b,b,b}=Reshape(bone_{n,b})$ (6)
$\Delta W_{n,m}=Reshape(W_{m/b,n/b,b,b}bone_{n/b,b,b}+bone_{n/b,b,b})$ (7)
In this way, we can easily adapt to any dimension of W and models with different architectures. The compatibility method used above groups by columns, and naturally, we can also group by rows. The visualization is shown in Figure  2(d), and the formula is as follows:
$W_{n/b,m/b,b,b}=Reshape(W_{n,m})$ (8)
$bone_{m/b,b,b}=Reshape(bone_{m,b})$ (9)
$\Delta W_{n,m}=Reshape(W_{n/b,m/b,b,b}bone_{m/b,b,b}+bone_{m/b,b,b})$ (10)
(a) LoRA
(b) Bone-both
(c) Bone-col
(d) Bone-row
@@ 4 Experiments
The experiments were conducted on 4×NVIDIA 4090 24G GPUs. Our experimental training parameter settings are shown in the table  1.
Table 1: Training parameter settings
Deivce Model low rank gobal batchsize learning rate precision
4×NVIDIA 4090 RWKV6 r=32/b=64 64 2e-5 bf16
Llama r=36/b=64 64 2e-5 bf16
4.1 Evaluating the Performance of Bone under Different Parameter Settings in RWKV6
@@
To validate the stability of Bone, we trained RWKV6 models with different parameter sizes on math tasks.
Math: We fine-tuned RWKV6 models with varying parameters on the MetaMathQA dataset Yu et al. (2023) and evaluated them on GSM8k Cobbe et al. (2021) and MATH datasets Yu et al. (2023).
@@
For fine-tuning RWKV6 models with different parameter sizes, we kept the learning rate, batch size, and training steps the same across all experiments. We set lora_r=32 and bone_b=64 to ensure that the trainable parameters are equivalent between the two methods.
Result
As shown in Table 2, Bone consistently outperforms PiSSA and LoRA across all parameter settings and tasks during fine-tuning. For example, on the GSM8k test, Bone achieved improvements of 1.82%, 0.61%, and 2.28% over PiSSA across different model sizes. This further demonstrates that the Bone architecture is adaptable to models with various parameter settings and consistently surpasses other fine-tuning structures.
(a) RWKV6 loss
Figure 3: The image displays the loss curves for RWKV6 with different parameter settings, fine-tuned on MetaMathQA.
Table 2: The performance of different PEFT techniques under various parameter settings on RWKV6.
Model Strategy Trainable GSM8K Math
RWKV 6-1.6B LoRA 33.0M 10.84 1.14
PiSSA 33.0M 12.89 1.28
Bone 33.0M 14.71 1.94
RWKV 6-3B LoRA 55.1M 21.22 2.34
PiSSA 55.1M 24.64 2.54
Bone 55.1M 25.25 2.82
RWKV 6-7B LoRA 88.1M 38.13 6.06
PiSSA 88.1M 40.48 6.12
Bone 88.1M 42.76 6.34
4.2 Experiments on Llama2-7B
@@
To verify the generalizability of Bone, we conducted more comprehensive experiments on LLaMA2-7B. we conducted 3 more task finetuning experiments on LLaMa2-7B: math, code, and chat.
@@
[…]
@@
The results, as shown in Table  3 and Figure  1(a), demonstrate that Bone outperforms other PEFT methods in terms of convergence speed, data fitting, and generalization capabilities. Specifically, Bone outperforms PISSA by 5.84% on GSM8K and by 4.89% on HumanEval, further Bone outperformed methods compared, on all 3 different tasks focusing on various disciplines, demonstrating the generalizability.
Table 3: We fine-tuned LLaMA2-7B using Bone and various LoRA variants, and evaluated performance on GSM8k, Math, HumanEval, and MT-Bench.
@@
Llama2-7B LoRA 89.9M 38.97 0.44 16.46 3.93
OLoRA 89.9M 41.93 6.48 20.12 4.09
PiSSA 89.9M 43.52 6.92 20.73 4.25
Bone 87.0M 49.36 8.88 25.62 4.81
4.3 Experiments on different values of block_size $(b)$.
This subsection explores the upper limits of the Bone structure by varying the block_size in the Bone matrix. Notably, while we change the size of $b$, the value of $n$ remains constant. Therefore, as $b$ increases, the dimensions of Bone transform into the bone_g form, leading to a corresponding reduction in the number of blocks within each group. Comparative experiments were conducted by fine-tuning LLaMA2-7B on the MetaMathQA dataset and validating on GSM8K and Math benchmarks. The test results, as shown in Table  4, demonstrate that the fine-tuning performance improves as the value of b increases. Notably, when b reaches 16, the Bone structure, with only one-quarter of the trainable parameters compared to PiSSA, surpasses PiSSA’s performance on the GSM8k benchmark. However, its performance on the Math benchmark is only 3.54. The GSM8K score surpasses that of PiSSA, but the Math score is significantly lower, indicating The size of b impacts the model’s ability to understand unseen data. Based on this observation, we conclude that increasing b not only raises the number of parameters but also enlarges the window for information fusion, allowing the model to perform better on unseen data.
Table 4: Comparing different values of block_size $(b)$

### v4

# Bone: Block Affine Adaptation of Large Language Models

@@ Abstract
Low-Rank Adaptation (LoRA) has achieved remarkable training results by freezing the original weights and training only low-rank matrices, establishing itself as the predominant fine-tuning method for LLMs. In pursuit of performance closer to full-parameter training, a series of LoRA variants have emerged, such as LoRA+, PISSA, Olora, and LoRA-GA. This paper introduces a novel PEFT technique distinct from LoRA, called Block-Affine Adaptation (Bone). By dividing the original weights into multiple subspaces that share a single matrix for weight updates, Bone simplifies the process by requiring the trainable matrix to be initialized to zero, eliminating the need for complex initialization as in some LoRA variants. Compared to LoRA, Bone significantly reduces memory usage and achieves faster computation. Evaluation of both NLU and NLG tasks demonstrates that Bone substantially outperforms LoRA and its variants. Inspired by Pissa, we further proposed the “Weight Guide” theory to better utilize the information from the original weights. By integrating “Weight Guide” with Bone, we developed a new structure called Block-Affine Transformation (Bat), and ablation experiments confirmed the effectiveness of “Weight Guide”.
@@ 1 Introduction
LoRA significantly reduces memory usage by freezing the original weights $W$ and updating two low-rank matrices $A$ and $B$. Typically, either $A$ or $B$ is initialized to zero, ensuring that the initial state of LoRA is consistent with the pre-trained model, The figure  2 illustrates the structure visualization. However, extensive experiments Ding et al. (2023); Liu et al. (2024b); Biderman et al. (2024) have shown that LoRA’s convergence is significantly slower compared to full fine-tuning. This slow convergence is likely due to the small gradients caused by the zero initialization of either $A$ or $B$. To address this issue, researchers have proposed several LoRA variants, such as LoRA+ Hayou et al. (2024), PISSA Meng et al. (2024), and LoRA-GAWang et al. (2024). Despite their excellent performance, these LoRA variants inevitably introduce complexity into the fine-tuning process and reduce structural flexibility. For example, LoRA+ requires manual adjustment of different learning rates for $A$ and $B$; PISSA necessitates performing SVD decomposition on $W$ at the beginning of training, which can be time-consuming when the model parameters are large. As research on LoRA becomes increasingly saturated, it is essential to explore new directions for PEFT techniques and effectively leverage the unique characteristics of LLMs.
In this work, we first designed Block-Affine Adaptation (Bone), which divides the original weights into multiple subspaces that share a single trainable matrix for updates, making it more efficient than LoRA’s two low-rank matrices. Secondly, to address the limitations of LoRA and fully utilize the information in the original weights, we propose a new theory, “Weight Guide”. “Weight Guide” enables feature extraction by applying a simple linear transformation to the original weights. Finally, to validate the effectiveness of “Weight Guide”, we combined Bone with “Weight Guide” to develop Block-Affine Transformation (Bat).
Our extensive evaluation shows that Bone not only excels in both Natural Language Understanding (NLU) and Natural Language Generation (NLG) tasks but also significantly outperforms LoRA and its variants. Additionally, Bone retains the advantages of LoRA, such as ease of use and no additional computational overhead for the model. In NLG tasks, Bone demonstrates superior performance, with evaluation metrics surpassing even the strong LoRA variant Pissa across the board. To verify the feasibility of Bone, we conducted experiments on two different LLM architectures (LLaMA2 Xu et al. (2023), RWKV6 Peng et al. (2024)). As shown in Figure  1, Bone achieves the fastest convergence, and the results on the test set demonstrate its superior data fitting and generalization capabilities ( Table  2). As a completely new structure distinct from LoRA, Bone offers clear improvements in computational efficiency and memory savings, as detailed in table  7.
@@
1. We propose a novel PEFT technique called Block-Affine Adaptation (Bone). Bone outperforms LoRA and its variants across various fine-tuning tasks. Additionally, Bone is more memory-efficient and computationally faster compared to LoRA.
2. To effectively leverage the implicit information in the original weights, we propose the “Weight Guide” theory. By integrating this theory with Bone, we design Block-Affine Transformation (Bat), demonstrating the effectiveness of “Weight Guide”.
@@ 2 Related Works
3 Method
3.1 Motivation
The primary motivations are as follows:
Firstly, current research on LoRA has reached a saturation point, underscoring the need for a new direction in Parameter-Efficient Fine-Tuning (PEFT) techniques.
Secondly, existing PEFT methods lack a thorough exploration and effective utilization of the original model weights.
3.2 Bone: Block-Affine update matrices
To design a novel PEFT structure that rivals LoRA, we conducted an in-depth study of dense matrices. Current theories suggest that LLM weights exhibit sparse properties. For example, LoRA utilizes low-rank matrices to update dense matrices, while techniques like GQA Ainslie et al. (2023) and MQA Shazeer (2019) employ shared mechanisms to reduce parameter overhead with remarkable effectiveness.
Inspired by these approaches, we hypothesize that dividing LLM weights into multiple subspaces allows different subspaces to share a single low-rank matrix for updates. As illustrated in the figure  2, we name this structure Block-Affine Adaptation (Bone).
It is evident that the ways to partition LLM weights are diverse, as illustrated in the figure. Different partitioning methods influence the shape of the Bone matrix, the computation logic, and ultimately, the processing speed. In this work, we propose the Best Bone structure. For a pre-trained weight matrix $W\in\mathbb{R}^{d\times k}$, we update it using a single matrix: $W+\Delta W=W+expand(B)$, where $B\in\mathbb{R}^{r\times d}$, and the rank $r\ll k$. This design not only retains the simplicity and ease of use characteristic of LoRA but also achieves higher computational efficiency. The calculation is expressed as follows:
$Y=Wx+\Delta Wx=Wx+sum(x)B$ (1)
(a) Bone instruciton
Figure 2: Bone requires training only a single low-rank matrix initialized to zero. Best Bone represents the optimal design structure of Bone. Subsequent experiments default to using Best Bone.
3.3 “Weight Guide”
It is well-known that LLMs have extremely large parameters, with complex interactions among their internal weights, but most LoRA variants utilize original weight information minimally, with only a few leveraging it during initialization to enhance performance. To increase the utilization of this information and promote internal interactions among weights, we propose a novel concept called “Weight Guide” This approach ensures that the trainable matrices are consistently guided and constrained by the original weights throughout every step of the training process. As illustrated in Figure  1(b), LoRA variants interact with the original weights only once during initialization to extract essential components. In contrast, the Bone structure is guided (constrained) by the original weights after each update, significantly increasing the new weights’ utilization of the original weights. As the training steps increase, the influence of the original weights becomes more pronounced. This is evident from the loss curves, which show that the differences between Bone and other methods grow larger in the later stages of training.
So, how can we extract the critical information from the original weights? A straightforward approach is to apply a learnable matrix directly to the original weights using the Hadamard product. This method essentially assigns a weight to each element, allowing the learnable matrix to autonomously determine the importance of individual elements. Although this method is simple and straightforward, the elements within the weights remain isolated, making it impossible to learn interactions between them. Therefore, we replace the Hadamard product with matrix multiplication, essentially applying a simple linear transformation. This approach links the entire weight matrix space, providing a better chance to discover the optimal solution. The corresponding formula is shown below:
$\Delta W_{n,n}$ $=W_{n,n}\odot bone_{n,n}$ (2)
$\Delta W_{n,n}$ $=W_{n,n}\otimes bone_{n,n}$
3.4 Bat: Block-Affine Transformation
In the previous section, we introduced Bone, which is both simple and efficient. However, applying the same updates to different subspaces of the weight matrix $W\in\mathbb{R}^{d\times k}$ is clearly counterintuitive. To allow for more flexibility in the update directions of different subspaces, we combine Bone with “Weight Guide”, proposing a new method called Block-Affine Transformation (Bat). To balance computational efficiency, we divide the original weights into smaller subspaces compared to Bone. Here, we name the trainable matrix $bone\in\mathbb{R}^{r\times d}$, where its rank $r$ is replaced with block size $b$. The formula for Bone-col is as follows:
$W_{k/b,d/b,b,b}$ $=Reshape(W_{d,k})$ (3)
$bone_{d/b,b,b}$ $=Reshape(bone_{b,d})$
$\Delta W_{d,k}$ $=Reshape(W_{k/b,d/b,b,b}\otimes bone_{d/b,b,b}+bone_{d/b,b,b})$
Similar to Bone, Bat can also adopt various grouping methods, as illustrated in the figure 3. In Section 6, we compared the performance of different partitioning strategies for the structure. The formula for Bone-row is as follows:
$W_{d/b,k/b,b,b}$ $=Reshape(W_{d,k})$ (4)
$bone_{k/b,b,b}$ $=Reshape(bone_{b,k})$
$\Delta W_{d,k}$ $=Reshape(W_{d/b,k/b,b,b}\otimes bone_{k/b,b,b}+bone_{k/b,b,b})$
(a) Bone-both
(b) Bone-col
(c) Bone-row
@@ 4 Experiments
In this section, we evaluate the performance of Bone on various benchmark datasets. Initially, we assess Natural Language Understanding (NLU) capabilities using a subset of the GLUE dataset with the robert-base model. Subsequently, we evaluated the Natural Language Generation (NLG) capabilities by fine-tuning the LLM.
The experiments were conducted on 4×NVIDIA 4090 24G GPUs.
4.1 Experiments on Natural Language Understanding
@@
We fine-tune the RoBERTa-base model on several datasets from the GLUE benchmark, including MNLI, SST-2, CoLA, QNLI, and MRPC. Performance is evaluated on the development set using accuracy as the primary metric.
@@
The experimental hyperparameter settings were aligned with those in the LoRA repository, but training was conducted using a single 4090 GPU. Each experiment is conducted with 3 different random seeds, and the average performance is reported.
Results
As shown in Table 1, Bone demonstrates outstanding performance, particularly on the CoLA dataset, where it exhibits significantly faster convergence and superior data-fitting capabilities, far surpassing LoRA and Pissa.
Table 1: The results of fine-tuning RoBERTa-base using Bone and various LoRA variants were compared on a subset of the GLUE benchmark.
Method Trainable MNLI SST-2 CoLA QNLI MRPC
LoRA 0.2365% $85.63_{\pm 0.01}$ $\mathbf{94.03_{\pm 0.02}}$ $62.40_{\pm 0.71}$ $91.37_{\pm 0.97}$ $87.98_{\pm 0.23}$
Pissa 0.2365% $\mathbf{85.72_{\pm 0.40}}$ $93.64_{\pm 0.13}$ $67.28_{\pm 0.59}$ $91.40_{\pm 0.54}$ $88.11_{\pm 0.24}$
Bone 0.2360% $85.71_{\pm 0.32}$ $93.60_{\pm 0.07}$ $\mathbf{72.86_{\pm 3.13}}$ $\mathbf{91.43_{\pm 0.76}}$ $\mathbf{88.14_{\pm 0.60}}$
4.2 Experiment on Natural Language Generation
@@
To verify the generalizability of Bone, we conducted more comprehensive experiments on LLM. we conducted 3 more task finetuning experiments on LLM: math, code, and chat.
@@
The hyperparameter settings for this experiment were kept equal, while the train steps were adjusted according to the specific fine-tuning datasets used. It is worth noting that the weights of LLaMA2-7B are not fully symmetric, making it impossible to perfectly align the trainable parameters when comparing Bone and LoRA. To address this, we set the rank $r$ of LoRA to 36 and the rank $r$ of Bone to 64, ensuring that Bone uses fewer parameters than LoRA to demonstrate its superiority. Each experiment is conducted with 2 different random seeds, and the average performance is reported.
@@
The results, as shown in Table  2 and Figure  1(a), demonstrate that Bone outperforms other PEFT methods in terms of convergence speed, data fitting, and generalization capabilities. Bone demonstrates outstanding performance across three different tasks. On LLaMA2-7B, Bone achieves results that surpass Pissa, despite using fewer parameters than LoRA and its variants. On RWKV6-7B, Bone and LoRA have the same number of trainable parameters, yet Bone consistently outperforms LoRA and its variants across all tasks.
Table 2: We fine-tuned LLMs using Bone and various LoRA variants, and evaluated performance on GSM8k, Math, HumanEval, and MT-Bench.
@@
Llama2-7B LoRA 89.9M 40.75 5.22 17.68 3.73
OLoRA 89.9M 42.93 6.51 21.12 4.03
PiSSA 89.9M 43.89 6.92 22.25 4.11
Bone 87.0M 48.16 8.58 24.08 4.31
RWKV 6-7B LoRA 88.1M 38.13 6.06 - -
PiSSA 88.1M 40.48 6.12 - -
Bone 88.1M 41.73 6.52 - -
4.3 Effect of Rank $r$
This subsection explores the upper limits of the Bone structure by varying the rank $r$ in the Bone matrix. Comparative experiments were conducted by fine-tuning LLaMA2-7B on the MetaMathQA dataset and validating on GSM8K and Math benchmarks. The test results, as shown in Table  3, demonstrate that the fine-tuning performance improves as the value of b increases. Notably, when $r=16$, the Bone structure, with only one-quarter of the trainable parameters compared to PiSSA, surpasses PiSSA’s performance on the GSM8k benchmark. However, its performance on the Math benchmark is only 3.73. The GSM8K score surpasses that of PiSSA, but the Math score is significantly lower, indicating The size of $r$ impacts the model’s ability to understand unseen data. Based on this observation, we hypothesize that when the rank is too small, it significantly limits the model’s generalization ability.
Table 3: Comparing different values of rank $(r)$
Model rank Trainable GSM8K Math
Llama2-7B 16 21.7M 45.90 3.77
32 43.5M 46.18 7.43
64 87.0M 48.16 8.58
128 174.0M 53.49 10.08
Figure 4: Training loss curves of Bone with different rank $r$ on the MetaMathQA dataset.
4.4 Bone vs Bat
To validate the effectiveness of the “Weight Guide”, we fine-tuned LLaMA2-7B and RWKV6-7B using both Bone and Bat on the MetaMathQA dataset and evaluated their performance on Math and GSM8K. As shown in the table, Bat, equipped with “Weight Guide”, achieves significant improvements in performance metrics compared to Bone.
Table 4: Comparing Bat-row, Bat-col on math tasks